# Vision LLMs Are Bad at Hierarchical Visual Understanding, and LLMs Are the Bottleneck

### v1

@@ 6 Conclusion
This work presents a systematic evaluation of state-of-the-art VLLMs’s hierarchical visual understanding performance. We find that both open-source VLLMs and the proprietary GPT-4o give rise to low hierarchical consistency over six taxonomies of visual concepts. Probing results reveal that the visual and text embeddings carry rich hierarchical and discriminative cues, whereas the LLMs fail to decode them, implying LLMs are the bottleneck. Finetuning on hierarchical VQA tasks improves VLLMs’ hierarchical consistency on visual inputs while preserving their performance on general VQA tasks. Intriguingly, the finetuning benefits the LLM’s (text) hierarchical consistency more than the corresponding VLLM’s (visual) hierarchical measure. Ingesting the taxonomy-knowledge gap to LLMs, likely during pretraining rather than post-hoc patching, is a promising path toward VLLMs that reason coherently across different levels of semantic granularity about the visual world.
Appendices
In the appendices, we provide all implementation details to promote reproducibility of our work, more experimental results, and further discussions about this work. Section A is about the hierarchical image and text classification datasets. Section B supplements the main experiments in the paper, including the models, experiment setups, quantitative and qualitative results, and ablation studies. Section C presents various prompting and linear probing results, including those with the Gemma model and larger VLLMs up to 72B parameters. Section C allows one to reproduce our finetuning results and shows comparison results of text-only finetuning and on general VQA benchmarks. Finally, Sections EF, and G broaden the discussion by including limitations, broader impacts of the work, and more related works.
@@ A.1 Hierarchical Image Classification Benchmarks
Following prior work on hierarchical image classification, we adopted several commonly used hierarchical classification datasets, including ImageNet [13], iNaturalist-2021 [53], CUB-200-2011 [54] and Food-101 [5]. Due to the inherent unconstrained nature of open-ended predictions by VLLMs, even when provided with detailed instructions, their performance in open-ended hierarchical classification remains extremely limited, with an $\mathrm{Acc_{leaf}}$ as low as 3.88% by Qwen2.5-VL-7B. To more effectively evaluate model performance, we construct approximately one million multiple-choice questions in a four-choice VQA format. We provide the data construction process of hierarchy VQA benchmarks shown in Figure 7. To better illustrate the data format, we also provide several examples from different datasets as shown in Figure 8.
Figure 7: Overview of hierarchical image classification benchmarks construction process. Our hierarchical VQA benchmark is built on four datasets and covers six taxonomies. We first obtain the hierarchical structure for each taxonomy (biology standard and WordNet  [37] semantics). Then, we use SigLIP [66] to generate four choices for each image based on the image-text similarities. Finally, we leverage GPT to generate the corresponding questions.
@@ B.1 Models
Model Source
LLaVA-OV-7B [29] https://huggingface.co/lmms-lab/llava-onevision-qwen2-7b-ov
InternVL2.5-8B [9] https://huggingface.co/OpenGVLab/InternVL2_5-8B
InternVL3-8B [72] https://huggingface.co/OpenGVLab/InternVL3-8B
Qwen2.5-VL-7B [4] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct
Qwen2.5-VL-32B [4] https://huggingface.co/Qwen/Qwen2.5-VL-32B-Instruct
Qwen2.5-VL-72B [4] https://huggingface.co/Qwen/Qwen2.5-VL-72B-Instruct
GPT-4o [^1] GPT-4o results reported in this paper use the gpt-4o-2024-04-01-preview model for image-based tasks and the gpt-4o-2024-11-20 model for text-only evaluations. [1] https://openai.com/api/
OpenCLIP [10] https://huggingface.co/laion/CLIP-ViT-L-14-laion2B-s32B-b82K
SigLIP [66] https://huggingface.co/google/siglip-so400m-patch14-384
@@ B.2 Hierarchical Evaluation Metrics
$\mathrm{S\text{-}POR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{L_{i}}\max_{1\leq a\leq b\leq L_{i}}\Bigl{[}(b-a+1)\,\prod_{j=a}^{b}\mathbbm{1}\bigl{[}f_{\theta}(x_{i};\mathcal{Y}_{j})=y^{i}_{j}\bigr{]}\Bigr{]}.$
This stricter definition penalizes sporadic correctness and encourages full-path consistency.
@@
A TOR value of 1 indicates that every neighbouring pair is correctly predicted, whereas lower scores reflect violations of pairwise hierarchical coherence.
@@ B.3 Evaluation Results with All Metrics
LLaVA-OV-7B [29] 58.46 42.06 35.01 83.56 70.56 72.36 63.36 26.33 44.74 55.50 43.08 37.09
InternVL2.5-8B [9] 66.58 55.10 47.34 84.59 76.08 74.71 65.65 35.35 45.96 57.82 43.20 39.48
InternVL3-8B [72] 69.80 53.62 51.75 86.34 79.33 77.72 62.96 31.35 42.48 62.54 48.83 44.72
Qwen2.5-VL-7B [4] 80.85 70.52 67.97 90.52 84.52 83.84 64.12 26.47 44.53 71.95 59.37 57.45
Qwen2.5-VL-32B [4] 86.86 81.62 78.71 92.14 87.89 87.00 69.25 38.82 50.55 76.14 65.37 63.90
Qwen2.5-VL-72B [4] 89.79 86.20 83.63 92.43 88.74 87.77 67.21 30.86 48.10 80.23 70.92 69.86
@@
OpenCLIP [10] 47.22 19.04 17.28 71.68 37.71 47.88 53.80 20.60 28.35 34.40 14.17 11.38
SigLIP [66] 66.56 45.84 41.79 78.95 48.46 59.24 50.90 16.15 24.77 34.67 16.97 15.00
@@
LLaVA-OV-7B [29] 26.47 4.53 60.31 45.96 45.53
InternVL2.5-8B [9] 27.65 8.52 66.26 57.07 53.50
InternVL3-8B [72] 35.40 11.93 69.00 59.13 55.55
Qwen2.5-VL-7B [4] 41.66 19.73 74.80 66.92 63.71
Qwen2.5-VL-32B [4] 26.90 46.98 78.38 72.09 68.93
Qwen2.5-VL-72B [4] 35.73 54.20 81.76 76.05 73.55
@@
OpenCLIP [10] 23.53 1.04 41.11 19.02 21.12
SigLIP [66] 12.71 2.15 38.24 38.24 33.95
@@ B.5 Results on CUB-200 and iNat21-Plant with Random Choices
LLaVA-OV-7B [29] 40.25 86.14 78.12 59.30 58.94
InternVL2.5-8B [9] 64.20 91.06 88.36 77.13 76.91
InternVL3-8B [72] 67.50 93.80 90.55 75.96 82.23
Qwen2.5-VL-7B [4] 82.34 97.15 95.05 87.78 90.23
@@
LLaVA-OV-7B [29] 28.41 69.04 75.36 58.53 60.03
InternVL2.5-8B [9] 36.45 75.97 80.25 60.81 67.22
InternVL3-8B [72] 51.94 89.70 87.19 70.96 76.28
Qwen2.5-VL-7B [4] 70.09 93.76 92.75 82.88 86.15
@@ B.7 Food-101 Results
A comprehensive evaluation on Food-101 with all metrics is shown in Table 11. On the Food-101 dataset, all models achieve relatively high fine-grained classification accuracy. Unlike other datasets, LLaVA-OV-7B attains the highest HCA on this benchmark among the 7B/8B open-source VLLMs, even though its leaf-level accuracy is not the highest.
B.8 Could the Poor Hierarchical Consistency Originate from the Four-choice VQA Format?
@@
LLaVA-OV-7B [29] 88.80 46.45 77.30 57.70 57.06
InternVL2.5-8B [9] 84.76 41.18 72.91 52.77 51.85
InternVL3-8B [72] 84.88 37.95 71.26 49.11 48.96
Qwen2.5-VL-7B [4] 90.51 43.11 75.15 53.48 53.13
Qwen2.5-VL-32B [4] 89.13 47.32 76.99 56.83 57.93
Qwen2.5-VL-72B [4] 92.02 52.00 80.46 60.95 62.33
@@
OpenCLIP [10] 93.89 37.53 72.73 49.76 44.83
SigLIP [66] 97.17 42.49 73.68 50.03 51.69
@@
GPT-4o [1] 95.67 55.60 82.97 63.03 66.79
@@ C.1.1 Prompt Variation
Across all benchmarks we employ five distinct prompt templates (Table 12), comprising both hierarchy-aware (Hierarchical) and general formulations (General). For CUB200, iNat21-Animal, and iNat21-Plant, we use two hierarchy-specific prompts and three general prompts. For ImgNet-Animal and ImgNet-Artifact, all five prompts are general because the corresponding taxonomy trees are highly unbalanced, making level-specific queries ill-posed. We report the results in Table 13, averaging performance separately over general (General Prompts) and hierarchy-aware prompts (Hierarchy Prompts). Overall, hierarchy-aware prompts outperform general prompts on CUB-200 and iNat21-Plant.
[…]
@@
LLaVA-OV-7B [29] General Prompts 11.44 43.44 35.58 65.45 3.88 27.24
@@
InternVL2.5-8B [9] General Prompts 19.81 45.58 37.20 65.12 4.98 28.11
@@
Qwen2.5-VL-7B [4] General Prompts 40.78 65.38 55.33 79.93 16.15 40.51

### v2

@@ 6 Conclusion
This work presents a systematic evaluation of state-of-the-art VLLMs’s hierarchical visual recognition performance. We find that both open-source VLLMs and the proprietary GPT-4o give rise to low hierarchical consistency over six taxonomies of visual concepts. Probing results reveal that the visual and text embeddings carry rich hierarchical and discriminative cues, whereas the LLMs fail to decode them, implying LLMs are the bottleneck. Finetuning on hierarchical VQA tasks improves VLLMs’ hierarchical consistency on visual inputs while preserving their performance on general VQA tasks. Intriguingly, the finetuning benefits the LLM’s (text) hierarchical consistency more than the corresponding VLLM’s (visual) hierarchical measure. Ingesting the taxonomy-knowledge gap to LLMs, likely during pretraining rather than post-hoc patching, is a promising path toward VLLMs that reason coherently across different levels of semantic granularity about the visual world.
Supplementary Material
In the appendices, we provide all implementation details to promote the reproducibility of our work, more experimental results, and further discussions. Section A is about the hierarchical image and text classification datasets. Section B supplements the main experiments in the paper, including the models, experiment setups, quantitative and qualitative results, and ablation studies. Section C presents various prompting and linear probing results, including those with the Gemma model and larger VLLMs up to 72B parameters. Section D allows one to reproduce our finetuning results and shows comparison results of text-only finetuning and on general VQA benchmarks. Finally, Sections E ,  F, and  G broaden the discussions by including more related works, limitations, and broader impacts of the work.
@@ A.1 Hierarchical Image Classification Benchmarks
Following prior work on hierarchical image classification, we adopted several commonly used hierarchical classification datasets, including ImageNet [14], iNaturalist-2021 [63], CUB-200-2011 [64] and Food-101 [7]. Table 5 summarizes the six taxonomies and four datasets we use to construct the VQA tasks.
Dataset #Levels #Leaf Nodes #Images Hierarchy Distribution
CUB-200 [64] 4 200 5,794 13-37-124-200
iNat21-Plant [63] 6 4,271 42.71K 5-14-85-286-1702-4271
iNat21-Animal [63] 6 5,388 53.88K 6-27-152-715-2988-5388
ImgNet-Animal [14] 11 397 19.85K 2-10-37-81-123-81-65-41-64-34-2
ImgNet-Artifact [14] 7 491 24.55K 5-40-147-204-162-62-44
Food-101 [7] 4 84 21.00K 6-29-40-24
Due to the inherent unconstrained nature of open-ended predictions by VLLMs, even when provided with detailed instructions, their performance in open-ended hierarchical classification remains extremely limited, with an $\mathrm{Acc_{leaf}}$ as low as 3.88% by Qwen2.5-VL-7B. To more effectively evaluate model performance, we construct approximately one million multiple-choice questions in a four-choice VQA format. We provide the data construction process of hierarchy VQA benchmarks shown in Figure 7. To better illustrate the data format, we also provide several examples from different datasets as shown in Figure 8.
Figure 7: Overview of hierarchical image classification benchmarks construction process. Our hierarchical VQA benchmark is built on four datasets and covers six taxonomies. We first obtain the hierarchical structure for each taxonomy (biology standard and WordNet  [43] semantics). Then, we use SigLIP [78] to generate four choices for each image based on the image-text similarities, comprising the groundtruth class name and the top three classes returned by SigLIP. Finally, we leverage GPT-4o to generate the corresponding questions.
A.2 Taxonomy Refinement
To refine the taxnonomy quality, we use Wikipedia and GPT-4o to carefully examine the hierarchical relations in each taxonomy.
@@ A.2 Taxonomy Refinement
Re-Verify the Taxonomies: For example, in the original WordNet hierarchy, “indigo bunting” is misclassified under “finch”. However, based on the established taxonomy, it belongs to the cardinal family. We corrected its hierarchical path to: animal → vertebrate → bird → oscine → cardinal → indigo bunting. We detected these errors using GPT-4o and then validated them using reliable taxonomies in Wikipedia.
Remove Ambiguous Paths: For example, in WordNet, “tusker” is assigned to the overly coarse path: animal → vertebrate → mammal → tusker. However, “tusker” is merely a colloquial term for an elephant, and WordNet already includes a more fine-grained and taxonomically accurate path for “elephant”: animal → vertebrate → mammal → placental → elephant → African elephant. We removed the “tusker” node, as it lacks specificity and overlaps with the existing, more precise elephant class.
Correct Subordinate Relationships Within the Same Hierarchy Level: In the ImgNet-Artifact dataset, the hierarchy provided by WordNet exhibits notable semantic inconsistencies, particularly where concepts at the same hierarchical level implicitly reflect subordinate relationships. For instance, under the category “device”, concepts such as “machine”, “instrument”, “musical instrument”, and “mechanism” are listed as siblings. However, “musical instrument” is a subtype of “instrument”, making the latter a hypernym of the former. Treating these as peers can lead to ambiguous or conflicting answers when classifying an image, as both may be considered valid even though one is semantically nested within the other. To resolve these issues, we used GPT-4o to analyze whether sibling category pairs exhibited valid hypernym-hyponym relationships systematically. We refined or removed problematic intermediate nodes and eliminated leaf nodes associated with overly coarse or semantically inconsistent hierarchy paths.
A.3 Hierarchical Text Classification Benchmarks
@@ B.1 Models
Model Source (Huggingface Repos and API Platform)
LLaVA-OV-7B [33] lmms-lab/llava-onevision-qwen2-7b-ov
LLaVA-OV-1.5-8B [2] lmms-lab/LLaVA-OneVision-1.5-8B-Instruct
InternVL2.5-8B [11] OpenGVLab/InternVL2_5-8B
InternVL3-8B [83] OpenGVLab/InternVL3-8B
Qwen2.5-VL-7B [6] Qwen/Qwen2.5-VL-7B-Instruct
Qwen2.5-VL-32B [6] Qwen/Qwen2.5-VL-32B-Instruct
Qwen2.5-VL-72B [6] Qwen/Qwen2.5-VL-72B-Instruct
Qwen3-VL-8B [5] Qwen/Qwen3-VL-8B-Instruct
Qwen3-VL-32B [5] Qwen/Qwen3-VL-32B-Instruct
Qwen3-Omni-30B-A3B [70] Qwen/Qwen3-Omni-30B-A3B-Instruct
GPT-4o [^1] GPT-4o results reported in this paper use the gpt-4o-2024-04-01-preview model for image-based tasks and the gpt-4o-2024-11-20 model for text-only evaluations. [47] OpenAI API
OpenCLIP [12] laion/CLIP-ViT-L-14-laion2B-s32B-b82K
SigLIP [78] google/siglip-so400m-patch14-384
BioCLIP [59] imageomics/bioclip
BioCLIP2 [22] imageomics/bioclip-2
@@ B.2 Hierarchical Evaluation Metrics
$\mathrm{S\text{-}POR}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{L_{i}}\max_{1\leq a\leq b\leq L_{i}}$
$\qquad\Bigl[(b-a+1)\prod_{j=a}^{b}\mathbbm{1}\bigl[f_{\theta}(x_{i};\mathcal{Y}_{j})=y^{i}_{j}\bigr]\Bigr].$ (4)
@@
A TOR value of 1 indicates that every neighboring pair is correctly predicted, whereas lower scores reflect violations of pairwise hierarchical coherence.
@@ B.3 Evaluation Results with All Metrics
LLaVA-OV-7B 58.46 42.06 35.01 83.56 70.56 72.36 63.36 26.33 44.74 55.50 43.08 37.09
LLaVA-OV-1.5-8B 69.26 55.64 50.74 87.22 80.77 79.55 65.26 29.72 46.48 60.70 49.76 43.68
InternVL2.5-8B 66.58 55.10 47.34 84.59 76.08 74.71 65.65 35.35 45.96 57.82 43.20 39.48
InternVL3-8B 69.80 53.62 51.75 86.34 79.33 77.72 62.96 31.35 42.48 62.54 48.83 44.72
Qwen2.5-VL-7B 80.85 70.52 67.97 90.52 84.52 83.84 64.12 26.47 44.53 71.95 59.37 57.45
Qwen2.5-VL-32B 86.86 81.62 78.71 92.14 87.89 87.00 69.25 38.82 50.55 76.14 65.37 63.90
Qwen2.5-VL-72B 89.79 86.20 83.63 92.43 88.74 87.77 67.21 30.86 48.10 80.23 70.92 69.86
Qwen3-VL-8B 80.25 72.99 68.35 90.69 85.04 84.50 64.29 27.42 45.50 72.96 63.06 60.15
Qwen3-VL-32B 86.49 82.32 78.55 91.30 87.19 86.09 68.31 37.30 49.71 72.08 62.27 59.06
Qwen3-Omni-30B-A3B 84.41 79.57 75.26 92.52 88.55 87.78 65.80 29.46 46.75 75.23 65.09 62.93
@@
OpenCLIP 47.22 19.04 17.28 71.68 37.71 47.88 53.80 20.60 28.35 34.40 14.17 11.38
SigLIP 66.56 45.84 41.79 78.95 48.46 59.24 50.90 16.15 24.77 34.67 16.97 15.00
BioCLIP 83.99 65.63 71.82 55.36 25.01 28.29 29.03 9.13 8.73 69.80 31.13 47.62
BioCLIP2 86.27 64.31 75.52 64.16 31.28 39.75 34.03 7.57 11.12 84.34 54.86 68.64
@@
LLaVA-OV-7B 26.47 4.53 60.31 45.96 45.53
LLaVA-OV-1.5-8B 27.08 8.42 66.35 58.71 53.57
InternVL2.5-8B 27.65 8.52 66.26 57.07 53.50
InternVL3-8B 35.40 11.93 69.00 59.13 55.55
Qwen2.5-VL-7B 41.66 19.73 74.80 66.92 63.71
Qwen2.5-VL-32B 46.98 26.90 78.38 72.09 68.93
Qwen2.5-VL-72B 54.20 35.73 81.76 76.05 73.55
Qwen3-VL-8B 39.28 19.54 75.39 69.08 65.23
Qwen3-VL-32B 51.74 32.44 80.69 74.70 72.08
Qwen3-Omni-30B-A3B 46.93 27.25 78.76 72.50 69.58
@@
OpenCLIP 23.53 1.04 41.11 19.02 21.12
SigLIP 12.71 2.15 38.24 38.24 33.95
BioCLIP 88.13 17.61 72.46 29.10 55.77
BioCLIP2 95.94 41.84 85.89 56.13 73.34
@@ B.5 Results on CUB-200 and iNat21-Plant with Random Choices
LLaVA-OV-7B 40.25 86.14 78.12 59.30 58.94
InternVL2.5-8B 64.20 91.06 88.36 77.13 76.91
InternVL3-8B 67.50 93.80 90.55 75.96 82.23
Qwen2.5-VL-7B 82.34 97.15 95.05 87.78 90.23
@@
LLaVA-OV-7B 28.41 69.04 75.36 58.53 60.03
InternVL2.5-8B 36.45 75.97 80.25 60.81 67.22
InternVL3-8B 51.94 89.70 87.19 70.96 76.28
Qwen2.5-VL-7B 70.09 93.76 92.75 82.88 86.15
@@ B.7 Food-101 Results
A comprehensive evaluation on Food-101 with all metrics is shown in Table 11. On the Food-101 dataset, all models achieve high fine-grained classification accuracy. Unlike other datasets, LLaVA-OV-7B attains the highest HCA on this benchmark among the 7B/8B open-source VLLMs, even though its leaf-level accuracy is not the highest.
@@
LLaVA-OV-7B 88.80 46.45 77.30 57.70 57.06
LLaVA-OV-1.5-8B 87.86 43.42 75.60 55.20 53.82
InternVL2.5-8B 84.76 41.18 72.91 52.77 51.85
InternVL3-8B 84.88 37.95 71.26 49.11 48.96
Qwen2.5-VL-7B 90.51 43.11 75.15 53.48 53.13
Qwen3-VL-8B 91.27 51.76 79.57 60.61 60.95
Qwen2.5-VL-32B 89.13 47.32 76.99 56.83 57.93
Qwen2.5-VL-72B 92.02 52.00 80.46 60.95 62.33
Qwen3-VL-8B 91.27 51.76 79.57 60.61 60.95
Qwen3-VL-32B 89.92 45.53 77.12 55.89 56.73
Qwen3-Omni-30B-A3B 94.71 46.81 78.10 55.38 58.22
@@
OpenCLIP 93.89 37.53 72.73 49.76 44.83
SigLIP 97.17 42.49 73.68 50.03 51.69
BioCLIP 27.82 2.48 26.29 10.81 7.25
BioCLIP2 75.89 14.10 45.19 20.81 18.39
@@
GPT-4o 95.67 55.60 82.97 63.03 66.79
B.8 HierarCaps Results
We further experiment with HierarCaps [1] under this work’s settings, and the results are presented in Table  12. Both CLIPs and VLLMs have poor hierarchical consistency and are especially worse on abstract and shorter captions at higher levels of the hierarchy (i.e., the first two levels). This is likely due to the noisy nature of the top-level labels in the HierarCaps dataset, where each image can be associated with multiple coarse-grained captions. For example, the image with the caption “A table with three plates with food and a man and a woman both holding utensil near plates” can reasonably be classified under both “persons” and “table” categories at the first layer of the hierarchy. We also compute the HCA for the last two levels (Level 3 and Level 4), which contain more concrete and longer captions. However, there remains a noticeable gap between the HCA and the leaf node accuracy (Level 4). In many cases, the model can correctly classify at level 4 while still struggling at level 3, suggesting that the model does not fully understand the hierarchical relationship across longer textual contexts.
Table 12: Hierarchical evaluation results on HierarCaps (Level Accuracy and HCA).
Model Level 4 Level 3 Level 2 Level 1 HCA (All) HCA (Last two levels)
OpenCLIP 68.80 50.70 28.40 17.10 5.70 45.30
SigLIP 72.90 52.50 26.80 15.50 5.70 49.20
LLaVA-OV-7B 78.60 60.70 32.60 22.20 9.10 55.60
Qwen2.5-VL-7B 77.10 58.40 30.50 21.10 7.10 54.00
B.9 Performance Analysis across Different Datasets
Among all datasets, CUB-200 exhibits the smallest gap between HCA and leaf-level accuracy, which can be attributed to its shallow hierarchy (only four levels vs. six levels in iNat21-Plant and iNat21-Animal), making the task relatively simple compared to other datasets. ImgNet-Animal and ImgNet-Artifact have the deepest hierarchies. However, their leaf nodes generally correspond to basic-level concepts, which makes the leaf-level classification easier for VLLMs, resulting in high leaf accuracy. ImgNet-Artifact is the most challenging dataset on which all models yield low hierarchical consistency scores, probably for two main reasons. (1) Unlike the well-defined biological hierarchies in iNat-21 and ImgNet-Animal, the intermediate nodes in WordNet, which were used to construct ImageNet, are relatively abstract and vague, making hierarchical discrimination difficult. (2) Unlike the animal images, images in the ImgNet-Artifact dataset often contain multiple objects of different classes. When queried about higher-level categories, the model may mistakenly associate the question with another non-central object in the scene.
B.10 Could the Poor Hierarchical Consistency Originate from the Four-choice VQA Format?
@@ C.1.1 Prompt Variation
Across all benchmarks we employ five distinct prompt templates (Table 13), comprising both hierarchy-aware (Hierarchical) and general formulations (General). For CUB-200, iNat21-Animal, and iNat21-Plant, we use two hierarchy-specific prompts and three general prompts. For ImgNet-Animal and ImgNet-Artifact, all five prompts are general because the corresponding taxonomy trees are highly unbalanced, making level-specific queries ill-posed. We report the results in Table 14, averaging performance separately over general (General Prompts) and hierarchy-aware prompts (Hierarchy Prompts). Overall, hierarchy-aware prompts outperform general prompts on CUB-200 and iNat21-Plant.
Table 13: Prompt templates used across datasets. Placeholders: (i) CUB-200: level $\in$ {order, family, genus, species}; (ii) iNat21: object $\in$ {animal, plant}, level $\in$ {kingdom, phylum, class, order, family, genus, species}; (iii) ImgNet: class $\in$ {animal, artifact}.
@@
LLaVA-OV-7B General Prompts 11.44 43.44 35.58 65.45 3.88 27.24
@@
InternVL2.5-8B General Prompts 19.81 45.58 37.20 65.12 4.98 28.11
@@
Qwen2.5-VL-7B General Prompts 40.78 65.38 55.33 79.93 16.15 40.51